# A comparative analysis for different finite element types in strain-gradient elasticity simulations performed on Firedrake and FEniCS

### v2

@@ 2   Method of solution in strain-gradient elasticity
where we understand Einstein summation convention over the repeated indices, $i,j,k,\dots\in[1,2,3]$ in three-dimensional space, $\rho_{0}$ is the mass density of the material, and $\bm{u}$ is the displacement, defined by difference between the current position, $\bm{x}$, and the reference position, $\bm{X}$. The kinetic energy term in Lagrangian density is omitted since the problem is static; interested readers may refer to [55] for further details. The first term, $w$, is the deformation energy density, which is a function of the linearized Green–Lagrange strain and its gradient, $w=\hat{w}(\bm{\varepsilon},\nabla\bm{\varepsilon})$. The last term in the Lagrangian density represents the energy density due to the volumetric effects, including a body force density, $\bm{f}$, and a supply term, $\bm{\ell}$, which is accounted for work through displacement gradient. The linearized strain tensor and its gradient are defined by
@@
[…]
@@
$w(u_{z},u^{\prime}_{z},u^{\prime\prime}_{z})=$ $\frac{c_{1}}{8}(u_{z}^{\prime})^{4}+c_{2}\Big{(}\frac{(u_{z}^{\prime})^{2}}{2}+\frac{(u_{z}^{\prime})^{4}}{4}\Big{)}$ (21)
$+\frac{c_{3}}{2r}\Big{(}(u_{z}^{\prime})^{2}u_{z}^{\prime\prime}+(u_{z}^{\prime}+2ru_{z}^{\prime\prime})\Big{)}+2c_{4}(u_{z}^{\prime})^{2}(u_{z}^{\prime\prime})^{2}$
$+\frac{c_{5}}{4r^{2}}\Big{(}(u_{z}^{\prime})^{2}(1+4r^{2}(u_{z}^{\prime\prime})^{2})+r^{2}(u_{z}^{\prime\prime})^{2}+4r(u_{z}^{\prime})^{3}u_{z}^{\prime\prime}+2ru_{z}^{\prime}u_{z}^{\prime\prime}+(u_{z}^{\prime})^{4}\Big{)}$
$+\frac{c_{6}}{2r^{2}}\Big{(}(u_{z}^{\prime})^{2}(2r^{2}(u_{z}^{\prime\prime})^{2}+1)+r^{2}(u_{z}^{\prime\prime})^{2}+(u_{z}^{\prime})^{4}\Big{)}$
$+\frac{c_{7}}{4r^{2}}\Big{(}(u_{z}^{\prime})^{2}(4r^{2}(u_{z}^{\prime\prime})^{2}+1)+r^{2}(u_{z}^{\prime\prime})^{2}+(u_{z}^{\prime})^{4}\Big{)}$
@@
$\frac{\partial w}{\partial u^{\prime}_{z}}$ $=\frac{\lambda}{2}\left((u_{z}^{\prime})^{3}\right)+\mu\left((u_{z}^{\prime})+(u_{z}^{\prime})^{3}\right)+\frac{c_{3}}{2z}\left(3(u_{z}^{\prime})^{2}u^{\prime\prime}+4z(u_{z}^{\prime})(u^{\prime\prime})^{2}\right)+2c_{4}\left(2(u_{z}^{\prime})(u^{\prime\prime})^{2}\right)$
$+\frac{c_{5}}{4z^{2}}\biggl{(}2(u_{z}^{\prime})+8z^{2}(u_{z}^{\prime})u^{\prime\prime}(u_{z}^{\prime})^{2}+12z(u_{z}^{\prime})^{2}u^{\prime\prime}(u_{z}^{\prime})+2zu^{\prime\prime}+4(u_{z}^{\prime})^{3}\biggr{)}$
$+\frac{c_{6}}{2z^{2}}\left(4z^{2}(u_{z}^{\prime})(u^{\prime\prime})^{2}+2(u_{z}^{\prime})+4(u_{z}^{\prime})^{3}\right)+\frac{c_{7}}{4z^{2}}\left(8z^{2}(u_{z}^{\prime})(u^{\prime\prime})^{2}+2(u_{z}^{\prime})+4(u_{z}^{\prime})^{3}\right),$
$\frac{\partial w}{\partial u^{\prime\prime}_{z}}$ $=\frac{c_{3}}{2z}\left((u_{z}^{\prime})^{3}+4zu^{\prime\prime}(u_{z}^{\prime})^{2}\right)+2c_{4}\left(2u^{\prime\prime}(u_{z}^{\prime})^{2}\right)$
$+\frac{c_{5}}{4z^{2}}\biggl{(}8z^{2}(u_{z}^{\prime})^{2}u^{\prime\prime}+2z^{2}u^{\prime\prime}\delta u_{z}^{\prime\prime}+4z(u_{z}^{\prime})^{3}+2z(u_{z}^{\prime})+4(u_{z}^{\prime})^{3}\biggr{)}$
$+\frac{c_{6}}{2z^{2}}\left(4z^{2}(u_{z}^{\prime})^{2}u^{\prime\prime}++2z^{2}u^{\prime\prime}\right)+\frac{c_{7}}{4z^{2}}\left(8z^{2}(u_{z}^{\prime})^{2}u^{\prime\prime}++2z^{2}u^{\prime\prime}\right).$
@@ 3   Numerical results and discussion
$c_{3}=c_{4}=\frac{l_{c}^{2}}{112}\lambda$
$c_{5}=c_{7}=\frac{l_{c}^{2}}{1120}(7\mu+3\lambda)$
$c_{6}=\frac{l_{c}^{2}}{1120}(7\mu-4\lambda)\ ,$
@@
[…]
@@
[…]
r0ptb+220ptRefer to caption Refer to caption
@@
Table 3: CPU time(Assemble $+$ Computation) performance of different finite element formulations

### v3

@@ 2  Method of solution in strain-gradient elasticity
[…]
@@
$\mathcal{A}=\int_{\Omega}\mathcal{L}\,d\Sigma+\int_{\partial\Omega}W_{s}\,d\Gamma+\int_{\partial\partial\Omega}W_{e}\,d\Pi$ (7)
@@
$w(u_{z},u^{\prime}_{z},u^{\prime\prime}_{z})=$ $\frac{\lambda}{8}(u_{z}^{\prime})^{4}+\mu\Big(\frac{(u_{z}^{\prime})^{2}}{2}+\frac{(u_{z}^{\prime})^{4}}{4}\Big)$ (21)
$+\frac{a_{1}}{2r}\Big((u_{z}^{\prime})^{2}u_{z}^{\prime\prime}+(u_{z}^{\prime}+2ru_{z}^{\prime\prime})\Big)+2a_{2}(u_{z}^{\prime})^{2}(u_{z}^{\prime\prime})^{2}$
$+\frac{a_{3}}{4r^{2}}\Big((u_{z}^{\prime})^{2}(1+4r^{2}(u_{z}^{\prime\prime})^{2})+r^{2}(u_{z}^{\prime\prime})^{2}+4r(u_{z}^{\prime})^{3}u_{z}^{\prime\prime}+2ru_{z}^{\prime}u_{z}^{\prime\prime}+(u_{z}^{\prime})^{4}\Big)$
$+\frac{a_{4}}{2r^{2}}\Big((u_{z}^{\prime})^{2}(2r^{2}(u_{z}^{\prime\prime})^{2}+1)+r^{2}(u_{z}^{\prime\prime})^{2}+(u_{z}^{\prime})^{4}\Big)$
$+\frac{a_{5}}{4r^{2}}\Big((u_{z}^{\prime})^{2}(4r^{2}(u_{z}^{\prime\prime})^{2}+1)+r^{2}(u_{z}^{\prime\prime})^{2}+(u_{z}^{\prime})^{4}\Big)$
@@
$\frac{\partial w}{\partial u^{\prime}_{z}}$ $=\frac{\lambda}{2}\left((u_{z}^{\prime})^{3}\right)+\mu\left((u_{z}^{\prime})+(u_{z}^{\prime})^{3}\right)+\frac{a_{1}}{2z}\left(3(u_{z}^{\prime})^{2}u^{\prime\prime}+4z(u_{z}^{\prime})(u^{\prime\prime})^{2}\right)+2a_{2}\left(2(u_{z}^{\prime})(u^{\prime\prime})^{2}\right)$
[…]
$+\frac{a_{4}}{2z^{2}}\left(4z^{2}(u_{z}^{\prime})(u^{\prime\prime})^{2}+2(u_{z}^{\prime})+4(u_{z}^{\prime})^{3}\right)+\frac{a_{5}}{4z^{2}}\left(8z^{2}(u_{z}^{\prime})(u^{\prime\prime})^{2}+2(u_{z}^{\prime})+4(u_{z}^{\prime})^{3}\right),$
$\frac{\partial w}{\partial u^{\prime\prime}_{z}}$ $=\frac{a_{1}}{2z}\left((u_{z}^{\prime})^{3}+4zu^{\prime\prime}(u_{z}^{\prime})^{2}\right)+2a_{2}\left(2u^{\prime\prime}(u_{z}^{\prime})^{2}\right)$
$+\frac{a_{3}}{4z^{2}}\biggl(8z^{2}(u_{z}^{\prime})^{2}u^{\prime\prime}+2z^{2}u^{\prime\prime}\delta u_{z}^{\prime\prime}+4z(u_{z}^{\prime})^{3}+2z(u_{z}^{\prime})+4(u_{z}^{\prime})^{3}\biggr)$
$+\frac{a_{4}}{2z^{2}}\left(4z^{2}(u_{z}^{\prime})^{2}u^{\prime\prime}++2z^{2}u^{\prime\prime}\right)+\frac{a_{5}}{4z^{2}}\left(8z^{2}(u_{z}^{\prime})^{2}u^{\prime\prime}++2z^{2}u^{\prime\prime}\right).$
@@ 3  Numerical results and discussion
$c_{3}=c_{4}=\frac{l^{2}}{112}\lambda$
$c_{5}=c_{7}=\frac{l^{2}}{1120}(7\mu+3\lambda)$
$c_{6}=\frac{l^{2}}{1120}(7\mu-4\lambda)\ ,$
@@
$\text{Error}=\int_{0}^{1}|u^{\text{ref}}-u^{\text{sim}}|\,dx\approx\sum_{i=0}^{N-2}\frac{1}{2}\left(|u^{\text{ref}}_{i}-u^{\text{sim}}_{i}|+|u^{\text{ref}}_{i+1}-u^{\text{sim}}_{i+1}|\right)\Delta x$ (25)
@@
$a_{1}=\frac{l^{2}}{56}\lambda$ (26)
$a_{2}=\frac{l^{2}}{224}\lambda$
$a_{3}=\frac{l^{2}}{560}(7\mu+3\lambda)$
$a_{4}=\frac{l^{2}}{1120}(7\mu-4\lambda)$
$a_{5}=\frac{l^{2}}{560}(7\mu+3\lambda)\ ,$
@@
           Refer to captionRefer to caption
@@
Table 3: CPU time(Assemble ++ Computation) performance of different finite element formulations